OpenStudy (anonymous):

whats the slope of this equation ? 3x-5y=45 Is it 3/5 -3/5 3 or 5 ?

OpenStudy (anonymous):

solve the equation for y. have y on the left side of the = sign. make sure there is no number to the left of y when you are finished solving it. once you have solved for y let me know and we can go from there.

OpenStudy (anonymous):

3x-5y = 45 -5y = 45-3x // subtract the 3x from both sides i have it solved, but give it a go, and post your process step by step. if you get stuck just let me know.

OpenStudy (anonymous):

ok, -5y = 45-3x // we found that -y = (45/5)-(3/5)x // I divided both sides by 5 y = -9 +(3/5)x // I divided both sides by -1 and simplified y = (3/5)x -9 // I just moved the (3/5)x and -9 around therefore, you can find the slope since y = mx+b where m is the slope if you are having trouble with this concept the khan academy is a great resource for learning math. khanacademy.org